OpenStudy (anonymous):

A double-slit interference experiment is done in a ripple tank. The slits are 4.90 cm apart, and a viewing screen is 1.56 m from the slits. The wave speed of the ripples in water is 0.027 m/s, and the frequency of the oscillator producing the ripples is 5.10 Hz. How far from the centerline of the screen will the first order maxima be found?

OpenStudy (anonymous):

Let's call the position on the screen away from the centreline \(y\), the distance from the slits to the screen \(D\) = 1.56 m, the separation of the slits \(d\) = 4.9 cm, and the angle of diffraction \(\theta\). The wavelength \(\lambda\) is the wavespeed divided (0.027 m/s) by the frequency (5.1 Hz). The difference in path length from each slit to a point \(y\) on the screen is \(d\sin\theta\) where \(y\) is related to \(theta\) by \[\tan\theta=\frac{y}{D}\] The condition for the first order maximum is that the difference in path length must be 1 wavelength. \[d\sin\theta=\lambda\] \(y\) may be worked out thus: \[y=D\tan\sin^{-1}\Big(\frac{\lambda}{d}\Big)\] or you could use the small-angle approximation \[y = \frac{\lambda D}{d}\] In case the LaTeX on this website doesn't render for you (it doesn't for me), I've attached my answer compiled as a PDF